OpenStudy (anonymous):

The expression cos^2X - cos2X is equivalent to what?

OpenStudy (anonymous):

you know that \[cos(2x)=cos^2(x)-sin^2(x)\] so \[cos^2(x)-cos(2x)=cos^2(x)-[cos^2(x)-sin^2(x)]=sin^2(x)\]